OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1). a. y + 3 = −5(x − 2) b. y − 2 = −5(x + 3) c. y + 3 = −(x − 2) d. y − 2 = −1/5(x + 3)

OpenStudy (anonymous):

somebody please HELP me

OpenStudy (anonymous):

\[y _{2} - y _{1} = m ( x _{2}-x_{1}) , replace the points \to x2 and x1 \] got it ?

OpenStudy (anonymous):

so which one is it

OpenStudy (anonymous):

find m = y2 - y1 / x2 - x1

OpenStudy (anonymous):

im not really good at math can you HELP me out please !

OpenStudy (anonymous):

\[m = \frac{ 1-2 }{ 2-(-3)} = \frac{ -1 }{ 5}\]

OpenStudy (anonymous):

D is correct answer

OpenStudy (anonymous):

thanks